An isotropic point source emits light at wavelength 470 nm, at the rate of 230 W. A light detector is positioned 410 m from the source. What is the maximum rate dB/dt at which the magnetic component of the light changes with time at the detector's location? The speed of light is c = 3 × 108 m/s, and μ0 = 4π × 10-7 H/m.
Answer #1

Given is:-

Power P= 230W

Wavelength X= 470nm

Distance T= 410m

Speed of light C= 3 x 108m/s

μη = 4πX 10-H/m

Now,

The maximum rate dB/dt at which the magnetic component of the light changes with time is given by

(\frac{dB}{dt})_{max} = \sqrt{\frac{2 \mu_o P}{4\pi c}}(\frac{2 \pi c }{\lambda r})

by plugging all the values we get

2(47 x 10-7)(230), 26(3 x 108) 47(3 x 108) (470 x 10-9)(410)

which gives us

GB max = 3.8303 x 10ºT/s